Hi!
I need to find the following focus points of the following parabola!
y=7*x^2
y=2*x^2-2x-2

Answer by rothauserc(4718)

y = 7x^2
:
Note that a = 7 and the vertex of this parabola is (0, 0) then, the focus is (0, 0+(1/4a) = (0, 1/28)
:
y = 2x^2 -2x -2 = 2(x - (1/2))^2 - (5/2)
:
Note a = 2, and the vertex is ( (1/2), (-5/2) ) then the focus is ( (1/2), (-5/2 + (1/8) ) = ( (1/2), (-19/8) )
: